OpenStudy (anonymous):

What must be added to 3x -7 to make x2 + 4x -1?

OpenStudy (perl):

(3x - 7) + ? = x^2 + 4x - 1

OpenStudy (perl):

lets call what needs to be added A (3x - 7) + A = x^2 + 4x - 1 then by algebra A = ( x^2 + 4x - 1) - (3x-7)

OpenStudy (anonymous):

thank you i have been have a hard time with my algebra and this is the last answer on my homework and i have been struggling with it for a while now :)